# An Integral Equation Arising from Time-Consistent Portfolio Selection

### v1

@@ Abstract
This paper discusses a nonlinear integral equation arising from a class of time-consistent portfolio selection problem. We propose a unified framework requiring minimal assumptions, such as right-continuity of market coefficients and square-integrability of the market price of risk. Our main contribution is proving the existence and uniqueness of the square-integrable solution for the integral equation under mild conditions. Illustrative applications include the mean-variance portfolio selection and the utility maximization with random risk aversion.
@@ 1 Introduction
A time-inconsistent preference usually leads to the time-inconsistency of the optimal strategy: the strategy that is optimal today may not remain optimal tomorrow. To address the issue of time inconsistency, the intra-personal equilibrium, as a time-consistent solution to the problems with time-inconsistent preferences, is originally introduced by Strotz (1955). Since a precise definition of continuous-time intra-personal equilibrium was first provided by Ekeland and Lazrak (2006), the continuous-time intra-personal equilibrium has been studied for various time-inconsistent preferences. For instance, the mean-variance (MV) portfolio selection problem is investigated by Basak and Chabakauri (2010), Hu, Jin, and Zhou (2012, 2017), Björk, Murgoci, and Zhou (2014), and Dai, Jin, Kou, and Xu (2021); the non-exponential discounting problem is examined by Ekeland and Lazrak (2006), Ekeland and Pirvu (2008), Hamaguchi (2021), and Mbodji and Pirvu (2023). For a broader discussion on continuous-time intra-personal equilibrium, see Yan and Yong (2019), He and Zhou (2022), Hernández and Possamaï (2023), and the references therein.
@@
The continuous-time intra-personal equilibria for a class of time-inconsistent preferences can be characterized by the solutions to a class of nonlinear integral (or differential) equations. Below, we provide three specific preferences to illustrate this. The first one is the MV preference with the risk aversion parameter being inversely proportional to the wealth level, which is proposed by Björk, Murgoci, and Zhou (2014). In a market with constant coefficients, the authors apply the Arzelà-Ascoli theorem to establish the existence and uniqueness of the solution to their integral equation (4.11), thereby confirm the existence and uniqueness of the equilibrium strategy among state-independent strategies. The second one is the preference with random risk aversion of Desmettre and Steffensen (2023), which addresses the problem of maximizing utility with random risk aversion. This preference is time-inconsistent due to the inclusion of nonlinear functions of expectation. The authors characterize the equilibrium strategies by an infinite-dimensional system of ODEs. Specifically, when the risk aversion has two possible values, this system reduces to a three-dimensional ODE system. However, the existence and uniqueness of the solution to this system is not provided there. The third one is the generalized disappointment aversion (GDA) preference, which is discussed in our recent work Liang, Wang, Xia, and Yuan (2024). Under GDA, the implicit definition of certainty equivalence introduces time inconsistency. We characterize the equilibrium strategies by a nonlinear integral equation.
@@
This paper presents a unified approach to discuss the integral equation arising from a large class of time-consistent portfolio selection problems with portfolio bounds. Under mild conditions, we prove that this equation has a unique square-integrable solution.
Our main contribution is a unified framework that minimizes assumptions on market parameters. Specifically, we require only that the market coefficients are right-continuous and the market price of risk is square-integrable. In contrast, Björk, Murgoci, and Zhou (2014) and Desmettre and Steffensen (2023) assume constant market coefficients: the method in Björk, Murgoci, and Zhou (2014) cannot handle right-continuous coefficients, while Desmettre and Steffensen (2023) does not establish the existence and uniqueness of the solution to the ODE system. Liang, Wang, Xia, and Yuan (2024) assume that the market coefficients are right-continuous and bounded. Moreover, our equation can be applied to study the time monotonicity of equilibrium strategies. For instance, in the random risk aversion case discussed in Desmettre and Steffensen (2023), we theoretically demonstrate that in a single-stock market with constant market coefficients, the equilibrium investment decreases over time, exhibiting the so-called horizon effect.
@@
The remainder of the paper is organized as follows. Section 2 investigates the existence and the uniqueness of the solution of a specific type of integral equation. Section 3 derives this integral equation from a class of time-consistent portfolio selection problem. Finally, Section 4 provides two illustrative applications.
@@ 2 An Integral equation
$\leq\ell_{1}|\lambda(t)|\left[\sqrt{v_{a^{(1)}-a^{(2)}}(t)}+\sqrt{v_{\lambda}(t)}\cdot\sqrt{v_{a^{(1)}-a^{(2)}}(t)}\right]$
$\leq\ell_{1}|\lambda(t)|\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})}.$
@@
$\begin{split}\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})}\leq\ell_{1}\sqrt{v_{\lambda}(t_{1})}\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})}\quad\forall\,t_{1}\in[\tau,T).\end{split}$ (2.4)
@@
$\sqrt{v_{a^{(1)}-a^{(2)}}(t_{2},s_{1})}\leq\ell_{1}\sqrt{v_{\lambda}(t_{2},s_{1})}\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{2},s_{1})}\quad\forall\,t_{2}\in[\tau,s_{1}).$
@@
For every $\tau\in[0,T)$, let
$\mathcal{S}_{\tau}=\left\{a\in L^{2}((\tau,T),\mathbb{R}^{d}):v_{a}(\tau)\leq 1\right\}.$
Obviously, $\mathcal{S}_{\tau}$ is a closed subset of $L^{2}((\tau,T),\mathbb{R}^{d})$. For every $a\in\mathcal{S}_{\tau}$, let
@@
First, we show that $\mathcal{T}$ is a map from $\mathcal{S}_{\tau}$ to $\mathcal{S}_{\tau}$ for every $\tau\in[0,T)$ that is sufficiently close to $T$. Indeed, for every $a\in\mathcal{S}_{\tau}$, we have
$\sqrt{v_{a}(t)}\leq 1\text{ and }\left|y_{a}(t)\right|\leq\sqrt{v_{\lambda}(0)}\quad\forall t\in[\tau,T).$
Then
@@
Assuming further that $h$ is I-bounded, we have the global existence of the solution, as the following theorem shows.
@@
Let $h$ be locally Lipschitz. In the case where $\lambda\in L^{\infty}((0,T),\mathbb{R}^{d})$, $h$ is bounded, and no portfolio bounds are imposed, we have established the global existence and uniqueness of the solution in $L^{\infty}((0,T),\mathbb{R}^{d})$; see Liang, Wang, Xia, and Yuan (2024, Theorem 3.5). In contrast, Theorem 2.7 only requires $\lambda\in L^{2}((0,T),\mathbb{R}^{d})$ and $h$ is I-bounded to guarantee the existence and uniqueness of the solution in $L^{2}((0,T),\mathbb{R}^{d})$.
3 Time-consistent portfolio selection
We derive Equation (2.1) from a class of time-consistent portfolio selection problem in this section. Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},0\leq t\leq T\}$ is the augmented natural filtration generated by a standard $d$-dimensional Brownian motion $\{B(t),0\leq t\leq T\}$. In addition, suppose $\mathcal{F}=\mathcal{F}_{T}$. The market consists of one risk-free asset (bond) and $d$ risky assets (stocks). For simplicity, we assume that the interest rate of the bond is zero. The dynamics of the stock price processes $S_{i}$, $i=1,\cdots,d$ are given by
@@ 3 Time-consistent portfolio selection
For any $m\geq 1$ and $\mathbb{S}\subset\mathbb{R}^{m}$, $L^{0}(\mathbb{F},\mathbb{S})$ is the space of $\mathbb{S}$-valued, $\mathbb{F}$-progressively measurable processes. For each $t\in[0,T]$, $p\in[1,\infty]$, $L^{p}(\mathcal{F}_{t},\mathbb{S})$ is the set of all $L^{p}$-integrable, $\mathbb{S}$-valued, and $\mathcal{F}_{t}$-measurable random variables. A trading strategy is a process $\pi=\{\pi_{t},t\in[0,T)\}\in L^{0}(\mathbb{F},\mathbb{R}^{d})$ such that $\int_{0}^{T}|\pi^{\top}_{t}\mu(t)|\mathrm{d}t+\int_{0}^{T}|\sigma^{\top}(t)\pi_{t}|^{2}\mathrm{d}t<\infty$ a.s., where $\pi_{t}$ represents the vector of portfolio weights determining the investment of wealth into the stocks at time $t$. The self-financing wealth process $\{W^{\pi}_{t},0\leq t\leq T\}$ associated with a trading strategy $\pi$ satisfies the following stochastic differential equation (SDE):
[…]
The time-$t$ performance of a trading strategy $\pi$ is $J(t,\pi)$. We assume that $J(t,\cdot)$ can be (equivalently) represented as a functional of the conditional distribution of ${W^{\pi}_{T}\over W^{\pi}_{t}}$ given $\mathcal{F}_{t}$. Here are some examples that will be further discussed in the next section. [^3] Other examples include the MV preference for the log returns of Dai, Jin, Kou, and Xu (2021), the constant-relative-risk-aversion (CRRA) betweenness preference discussed by liang2023dynamic, and the GDA preference discussed by Liang, Wang, Xia, and Yuan (2024).
Example 3.1.
$\mathbb{E}_{t}\left[W^{\pi}_{T}\right]-\frac{\gamma}{2W^{\pi}_{t}}\textup{Var}_{t}\left[W^{\pi}_{T}\right],$ (3.2)
where $\gamma>0$ is a risk aversion index. The performance functional (3.2) is obviously equivalent to
$J(t,\pi)=\mathbb{E}_{t}\left[{W^{\pi}_{T}\over W^{\pi}_{t}}\right]-\frac{\gamma}{2}\textup{Var}_{t}\left[{W^{\pi}_{T}\over W^{\pi}_{t}}\right].$ (3.3)
Example 3.2.
$\int_{(0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{\pi}_{T}\right)\right]\right)\mathrm{d}\Gamma(\gamma),$ (3.4)
where $\Gamma$ is a distribution function on $(0,\infty)$ and $u^{\gamma}$ are the CRRA utility functions given by
$u^{\gamma}(x)=\begin{cases}\frac{x^{1-\gamma}}{1-\gamma},\quad\gamma>0,\gamma\neq 1,x>0,\\ \log x,\quad\gamma=1,x>0.\end{cases}$
The performance functional (3.4) is obviously equivalent to
$J(t,\pi)=\int_{(0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left({W^{\pi}_{T}\over W^{\pi}_{t}}\right)\right]\right)\mathrm{d}\Gamma(\gamma).$ (3.5)
@@
$\frac{W^{\bar{\pi}}_{T}}{W^{\bar{\pi}}_{t}}\sim\mathrm{LogNormal}\left(y_{a}(t)-{\frac{1}{2}}v_{a}(t),v_{a}(t)\right).$ (3.8)
As $J(t,\cdot)$ is a functional of the conditional distribution of ${W^{\pi}_{T}\over W^{\pi}_{t}}$ given $\mathcal{F}_{t}$, we have [^4] Here, we implicitly assume that $\mathcal{D}\subset\Pi$.$J(t,\bar{\pi})=g(t,v_{a}(t),y_{a}(t))$ for some $g:[0,T)\times[0,\infty)\times\mathbb{R}\to\mathbb{R}$.
@@
$a(t)=\mathcal{P}_{t}\left(-\frac{g_{y}(t,v_{a}(t),y_{a}(t))}{2g_{v}(t,v_{a}(t),y_{a}(t))}\lambda(t)\right),\quad t\in[0,T),$ (3.9)
@@
Then Equation (3.9) is exactly Equation (2.1). So we finish deriving Equation (2.1).

### v2

# An Integral Equation in Portfolio Selection with Time-Inconsistent Preferences

@@ Abstract
This paper discusses a nonlinear integral equation arising from portfolio selection with a class of time-inconsistent preferences. We propose a unified framework requiring minimal assumptions, such as right-continuity of market coefficients and square-integrability of the market price of risk. Our main contribution is proving the existence and uniqueness of the square-integrable solution for the integral equation under mild conditions. Illustrative applications include the mean-variance portfolio selection and the utility maximization with random risk aversion.
@@ 1 Introduction
A time-inconsistent preference usually leads to the time-inconsistency of the optimal strategy: the strategy that is optimal today may not remain optimal tomorrow. To address the issue of time inconsistency, the intra-personal equilibrium is originally introduced by Strotz (1955). Since a precise definition of continuous-time intra-personal equilibrium was first provided by Ekeland and Lazrak (2006), the continuous-time intra-personal equilibrium has been studied for various time-inconsistent preferences. For instance, the mean-variance (MV) portfolio selection problem is investigated by Basak and Chabakauri (2010), Hu, Jin, and Zhou (2012, 2017), Björk, Murgoci, and Zhou (2014), and Dai, Jin, Kou, and Xu (2021); the non-exponential discounting problem is examined by
Ekeland and Lazrak (2006), Ekeland and Pirvu (2008), Hamaguchi (2021), and Mbodji and Pirvu (2023). For a broader discussion on continuous-time intra-personal equilibrium, see Yan and Yong (2019), He and Jiang (2021), Hernández and Possamaï (2023), and the references therein.
@@
The continuous-time intra-personal equilibria for a class of time-inconsistent preferences can be characterized by the solutions to a class of nonlinear integral (or differential) equations. Below, we provide four specific preferences to illustrate this. The first one is the MV preference with the risk aversion parameter being inversely proportional to the wealth level, which is proposed by Björk, Murgoci, and Zhou (2014). In a market with constant coefficients, the authors apply the Arzelà-Ascoli theorem to establish the existence and uniqueness of the solution to their integral equation (4.11), thereby confirm the existence and uniqueness of the equilibrium strategy among state-independent strategies. The second one is the preference with random risk aversion of Desmettre and Steffensen (2023), which addresses the problem of maximizing utility with random risk aversion. This preference is time-inconsistent due to the inclusion of nonlinear functions of expectation. The authors characterize the equilibrium strategies by an infinite-dimensional system of ordinary differential equations (ODEs). Specifically, when the risk aversion has two possible values, this system reduces to a three-dimensional ODE system. However, the existence and uniqueness of the solution to this system is not provided there. The third one is the generalized disappointment aversion (GDA) preference, which is discussed in Liang, Wang, Xia, and Yuan (2024). Under GDA, the implicit definition of certainty equivalence introduces time inconsistency. The equilibrium strategies are characterized by a nonlinear integral equation. The fourth one is the (endogenous habit, or mean) scaled MV preferences and (endogenous habit formation, or generalized) mean–standard-deviation preferences of Kryger, Nordfang, and Steffensen (2020). The equilibrium strategy is characterized by their ODE system $(21)-(25)$ (in the case of endogenous habit, scaled MV preference, the existence and uniqueness of the solution to the ODE system remains unproved), which implies that the equilibrium strategy is absolutely continuous.
@@
This paper presents a unified approach to discuss the integral equation arising from portfolio selection with a large class of time-inconsistent preferences and portfolio bounds. We establish the existence and the uniqueness of solution to this equation.
Our main contribution is a unified framework that minimizes assumptions on market parameters. Specifically, we require only that the market coefficients are right-continuous and the market price of risk is square-integrable. In contrast, Björk, Murgoci, and Zhou (2014),
Kryger, Nordfang, and Steffensen (2020), and Desmettre and Steffensen (2023) assume constant market coefficients: the methods in Björk, Murgoci, and Zhou (2014) and Kryger, Nordfang, and Steffensen (2020) cannot handle right-continuous coefficients, while Desmettre and Steffensen (2023) and a part of Kryger, Nordfang, and Steffensen (2020)) do not establish the existence and uniqueness of the solution to their ODE system; Liang, Wang, Xia, and Yuan (2024) assume that the market coefficients are right-continuous and bounded.
@@
The remainder of the paper is organized as follows. Section 2 investigates the existence and the uniqueness of the solution of a specific type of integral equation. Section 3 derives this integral equation from portfolio selection with a class of time-inconsistent preferences. Finally, Section 4 provides two illustrative applications.
@@ 2 An Integral equation
$\leq$ $\ell_{1}|\lambda(t)|\left[\sqrt{v_{a^{(1)}-a^{(2)}}(t)}+\sqrt{v_{\lambda}(t)}\cdot\sqrt{v_{a^{(1)}-a^{(2)}}(t)}\right]\leq\ell_{1}|\lambda(t)|\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})}.$
@@
$\begin{split}\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})}\leq\ell_{1}\sqrt{v_{\lambda}(t_{1})}\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})},\quad\forall\,t_{1}\in[\tau,T).\end{split}$ (2.4)
@@
$\sqrt{v_{a^{(1)}-a^{(2)}}(t_{2},s_{1})}\leq\ell_{1}\sqrt{v_{\lambda}(t_{2},s_{1})}\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{2},s_{1})},\quad\forall\,t_{2}\in[\tau,s_{1}).$
@@
For every $\tau\in[0,T)$, let $\mathcal{S}_{\tau}=\left\{a\in L^{2}((\tau,T),\mathbb{R}^{d}):v_{a}(\tau)\leq 1\right\}$. Obviously, $\mathcal{S}_{\tau}$ is a closed subset of $L^{2}((\tau,T),\mathbb{R}^{d})$. For every $a\in\mathcal{S}_{\tau}$, let
@@
First, we show that $\mathcal{T}$ is a map from $\mathcal{S}_{\tau}$ to $\mathcal{S}_{\tau}$ for every $\tau\in[0,T)$ that is sufficiently close to $T$. Indeed, for every $a\in\mathcal{S}_{\tau}$, we have $\sqrt{v_{a}(t)}\leq 1\text{ and }\left|y_{a}(t)\right|\leq\sqrt{v_{\lambda}(0)}$, $\forall t\in[\tau,T)$. Then
@@
Assuming further that $h$ is I-bounded, we have the following global existence of the solution:
@@
Let $h$ be locally Lipschitz. In the case where $\lambda\in L^{\infty}((0,T),\mathbb{R}^{d})$, $h$ is bounded, and no portfolio bounds are imposed, we, with Yuan, have established the global existence and uniqueness of the solution in $L^{\infty}((0,T),\mathbb{R}^{d})$; see Liang, Wang, Xia, and Yuan (2024, Theorem 3.5). In contrast, Theorem 2.7 only requires $\lambda\in L^{2}((0,T),\mathbb{R}^{d})$ and $h$ is I-bounded to guarantee the existence and uniqueness of the solution in $L^{2}((0,T),\mathbb{R}^{d})$. The I-boundedness of $h$ can be guaranteed by the boundedness of $h$ or the boundedness of $\{\mathbb{U}_{t},t\in[0,1)\}$, which are easy to verify.
3 Portfolio selection with time-inconsistent preferences
We derive Equation (2.1) from portfolio selection with a class of time-inconsistent preferences in this section. Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},0\leq t\leq T\}$ is the augmented natural filtration generated by a standard $d$-dimensional Brownian motion $\{B(t),0\leq t\leq T\}$. In addition, suppose $\mathcal{F}=\mathcal{F}_{T}$. The market consists of one risk-free asset (bond) and $d$ risky assets (stocks). For simplicity, we assume that the interest rate of the bond is zero. The dynamics of the stock price processes $S_{i}$, $i=1,\cdots,d$ are given by
@@ 3 Portfolio selection with time-inconsistent preferences
For any $m\geq 1$ and $\mathbb{S}\subset\mathbb{R}^{m}$, $L^{0}(\mathbb{F},\mathbb{S})$ is the space of $\mathbb{S}$-valued, $\mathbb{F}$-progressively measurable processes. For each $t\in[0,T]$, $p\in[1,\infty]$, $L^{p}(\mathcal{F}_{t},\mathbb{S})$ is the set of all $L^{p}$-integrable, $\mathbb{S}$-valued, and $\mathcal{F}_{t}$-measurable random variables. A trading strategy is a process $\pi=\{\pi_{t},t\in[0,T)\}\in L^{0}(\mathbb{F},\mathbb{R}^{d})$ such that $\int_{0}^{T}|\pi^{\top}_{t}\mu(t)|\mathrm{d}t+\int_{0}^{T}|\sigma^{\top}(t)\pi_{t}|^{2}\mathrm{d}t<\infty$ a.s., where $\pi_{t}$ represents the vector of portfolio weights determining the investment of wealth into the stocks at time $t$. The self-financing wealth process $\{W^{\pi}_{t},0\leq t\leq T\}$ associated with a trading strategy $\pi$ satisfying the following stochastic differential equation (SDE):
$\mathrm{d}W^{\pi}_{t}=W^{\pi}_{t}\pi^{\top}_{t}\mu(t)\mathrm{d}t+W^{\pi}_{t}\pi^{\top}_{t}\sigma(t)\mathrm{d}B(t),\quad W^{\pi}_{0}=w_{0}>0.$ (3.1)
The time-$t$ performance of a trading strategy $\pi$ is $J(t,\pi)$. We assume that $J(t,\cdot)$ can be (equivalently) represented as a functional of the conditional distribution of ${W^{\pi}_{T}\over W^{\pi}_{t}}$ given $\mathcal{F}_{t}$. Some examples will be presented after Definition 3.1.
@@
Here are two examples of preference functionals $J(t,\cdot)$ that depend only on the conditional distribution of ${W^{\pi}_{T}\over W^{\pi}_{t}}$ given $\mathcal{F}_{t}$, which will be further discussed in the next section. [^3] Other examples include the (endogenous habit, or mean) scaled MV preferences and (endogenous habit formation, or generalized) mean–standard-deviation preferences of Kryger, Nordfang, and Steffensen (2020), the MV preference for the log returns of Dai, Jin, Kou, and Xu (2021), the constant-relative-risk-aversion (CRRA) betweenness preference discussed by Liang, Xia, and Yuan (2023), and the GDA preference discussed by Liang, Wang, Xia, and Yuan (2024).
Example 3.2.
$J_{0}(t,\pi)=\mathbb{E}_{t}\left[W^{\pi}_{T}\right]-\frac{\gamma}{2W^{\pi}_{t}}\textup{Var}_{t}\left[W^{\pi}_{T}\right]=W^{\pi}_{t}J(t,\pi),$ (3.3)
where $\gamma>0$ is a risk aversion index and
$J(t,\pi)=\mathbb{E}_{t}\left[{W^{\pi}_{T}/W^{\pi}_{t}}\right]-\frac{\gamma}{2}\textup{Var}_{t}\left[{W^{\pi}_{T}/W^{\pi}_{t}}\right].$ (3.4)
Then, under Definition 3.1, an equilibrium strategy for (3.3) is an equilibrium strategy for (3.4), and vice versa.
Example 3.3.
$J_{0}(t,\pi)=\int_{(0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{\pi}_{T}\right)\right]\right)\mathrm{d}\Gamma(\gamma)=W^{\pi}_{t}J(t,\pi),$ (3.5)
where $\Gamma$ is a distribution function on $(0,\infty)$, $u^{\gamma}$ are the CRRA utility functions given by
$u^{\gamma}(x)=\begin{cases}\frac{x^{1-\gamma}}{1-\gamma},\quad\gamma>0,\gamma\neq 1,x>0,\\ \log x,\quad\gamma=1,x>0,\end{cases}$
and
$J(t,\pi)=\int_{(0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left({W^{\pi}_{T}/W^{\pi}_{t}}\right)\right]\right)\mathrm{d}\Gamma(\gamma).$ (3.6)
Then, under Definition 3.1, an equilibrium strategy for (3.5) is an equilibrium strategy for (3.6), and vice versa.
@@
${W^{\bar{\pi}}_{T}}/{W^{\bar{\pi}}_{t}}=\mathrm{e}^{\int_{t}^{T}a^{\top}(s)\lambda(s)\mathrm{d}s-\frac{1}{2}\int_{t}^{T}|a(s)|^{2}\mathrm{d}s+\int_{t}^{T}a^{\top}(s)\mathrm{d}B(s)}=\mathrm{e}^{y_{a}(t)-\frac{1}{2}v_{a}(t)+\int_{t}^{T}a^{\top}(s)\mathrm{d}B(s)}.$
Therefore,
${\text{given the condition $\mathcal{F}_{t}$,}}\quad{W^{\bar{\pi}}_{T}}/{W^{\bar{\pi}}_{t}}\sim\mathrm{LogNormal}\left(y_{a}(t)-{\frac{1}{2}}v_{a}(t),v_{a}(t)\right).$ (3.8)
As $J(t,\cdot)$ can be represented as a functional of the conditional distribution of ${W^{\pi}_{T}\over W^{\pi}_{t}}$ given $\mathcal{F}_{t}$, we have [^4] Here, we implicitly assume that $\mathcal{D}\subset\Pi$.$J(t,\bar{\pi})=g(t,v_{a}(t),y_{a}(t))$ for some $g:[0,T)\times[0,\infty)\times\mathbb{R}\to\mathbb{R}$. Using this function $g$, we can characterize the equilibrium strategies by an integral equation of $a$, as shown in the following theorem.
@@
$a(t)=\mathcal{P}_{t}\left(-\frac{g_{y}(t,v_{a}(t),y_{a}(t))}{2g_{v}(t,v_{a}(t),y_{a}(t))}\lambda(t)\right),\quad t\in(0,T),$ (3.9)
@@
Then Equation (3.9) is exactly Equation (2.1). Thus we complete deriving Equation (2.1).